Solve the system of equations. (If the system is dependent, enter a general solution in terms of c
{2x− y+ z = 12
2y− 3z= −16
3y + 2z= 2
(x, y, z) =

Answer by ankor@dixie-net.com(22740)

Solve the system of equations. (If the system is dependent, enter a general solution in terms of c
2x− y+ z = 12
2y − 3z = −16
3y + 2z = 2
(x, y, z) =
:
Let's just deal with the last two equations
2y - 3z = -16
3y + 2z = 2
Use elimination, multiply the 1st eq by 2, the 2nd eq by 3
4y - 6z = -32
9y + 6z = 6
--------------addition eliminates z, find y
13y = -26
y = -26/13
y = -2
Find z using the 3rd equation
3(-2) + 2z = 2
-6 + 2z = 2
2z = 2 + 6
2z = 8
z = 8/2
z = 4
Find x using first original equation
2x - y + z = 12
y=-2, z=4
2x - (-2) + 4 = 12
2x + 2 + 4 = 12
2x = 12 - 6
2x = 6
x = 3
therefore we have x=3, y=-2, z= 4